# The fundamental localization phases in quasiperiodic systems: A unified framework and exact results

### v4

@@ IV.3 Unified picture for the seven fundamental localization phases
Reducing the chiral parameter $\eta$ from 1 to 1/2 corresponds to turning on the on-site potential $V_{0}$ in the type-II QP mosaic model, transitioning from $V_{0}=0$ to $V_{0}=t$. Breaking the chiral symmetry of both models, introducing the MEs between localized and critical states, leading to the L+C phase. As the on-site potential increases towards the high-symmetry line $V_{0}=t$, or when the QP optical Raman lattice model reaches $\eta=1/2$, the system becomes analytically solvable again. The type-II QP mosaic model exhibits analytic MEs $E=\pm\lambda$ separating the localized and critical states Zhou et al. (2023), which corresponds to the MEs $E=\pm 2t_{0}$ in the QP optical Raman lattice model.
@@ V Schemes for experimental realization
Finally, we propose experimental schemes for realizing the SSQP on-site model [Eq. (34)] and QP optical Raman lattice model [Eq. (38)]. These schemes utilize ultracold alkali atoms, which are ideal candidates due to their large fine-structure energy splitting and moderate natural linewidth. These properties facilitate efficient Raman coupling and spin-dependent potentials, even when the laser frequency is far detuned from the $D_{1}$ and $D_{2}$ lines Soltan-Panahi et al. (2011); Kuzmenko et al. (2019); Szulim et al. (2022), making ultracold alkali atoms well-suited for implementing the QP lattice.
@@
where the first term represents the kinetic energy. ${\cal V}_{p}(z)=V_{p}\cos(k_{p}z+\phi_{p})$ is a deep spin-dependent primary lattice that freezes the spin-conserved hopping for spin-down atoms. The different terms in Eq. (34) can be realized as follows. The uniform spin-conserved hopping $\Pi_{j}=t\Lambda_{+}$ for spin-up atoms can be obtained by
@@
where $w_{s,\uparrow}(z)$ is the Wannier function of the secondary lattice generated by spin-independent potential ${\cal V}_{s}(z)=V_{s}\cos(k_{s}z+\phi_{s})$, with $a_{s}=\pi/k_{s}$ being the lattice constant. The potential ${\cal V}_{s}(z)$ is incommensurate with the potential ${\cal V}_{p}(z)$ for spin-down atoms, leading to the QP potential $M_{j}=2V_{B}V_{j}^{\rm d}\Lambda_{-}$ for spin-down atoms
[…]
The proposed Hamiltonian in Eq. (57) can be realized based on QP optical Raman lattice, as outlined below. The detailed implementation can be found in Appendix. G. To facilitate the description, we perform a unitary transformation, such that $\sigma_{z}\rightarrow\sigma_{x}$ and $\sigma_{x}\rightarrow-\sigma_{z}$. The kinetic term and spin-independent potential ${\cal V}_{s}(z)$ remain unchanged and can be generated using the standing wave ${\bf E}_{1}$. The rotated spin-dependent potential ${\cal V}_{p}(z)(\sigma_{0}-\sigma_{x})/2$ can be realized using a standing waves ${\bf E}_{2}$ and ${\bf E}_{3}$ along with a traveling wave ${\bf E}_{4}$. Finally, the $M_{0}\sigma_{z}$ term arises from the two-photon detuning of the Raman coupling. After performing the inverse unitary transformation, we reach the Hamiltonian in Eq. (57). More details can be found in Appendix. G, where ${}^{87}{\rm Rb}$ atoms are used as an example.
@@
The different terms in Eq. (38) can be realized accordingly. The uniform hopping coupling $\Pi_{j}=t_{0}\sigma_{z}+it_{\rm so}\sigma_{y}$ arises from the tunneling of the Wannier function of the primary lattice, where
@@
with $p_{z}^{2}/2m$ being the kinetic energy along the lattice direction $z$, $a_{0}$ the primary lattice constant, and ${\cal V}_{p}(z)=V_{p}\cos^{2}(k_{p}z)$ the spin-independent primary lattice realized by a blue-detuned standing wave ${\bf E}_{2}$. The $t_{\rm so}$ is generated by the Raman potential
[…]
where ${\cal M}(z)=M_{0}\cos(k_{p}z)$ represents the Raman potential formed by the intersection of $x$-polarized ${\bf E}_{2}$ and $z$-polarized traveling wave ${\bf E}_{3}$, generating the spin-flipped hopping. The QP on-site matrix element is realized by the spin-dependent secondary lattice, which is given by
[…]
where the QP parameter $\alpha=k_{s}/k_{p}$, and ${\cal V}_{s}(z)=V_{s}\cos^{2}(k_{s}z+\phi_{s})$ is the spin-dependent secondary lattice generated by the composite standing wave ${\bf E}_{1}$, formed by two counter-propagating beams with mutually perpendicular polarization along the $x$ and $y$ axes. The chiral parameter $\eta$ is tuned by adjusting the power ratio of the two different polarizations of beam ${\bf E}_{1}$. The spin-dependent term $\eta{\cal V}_{s}(z)\sigma_{z}$ from vector optical shift is induced by the y-polarization, while the spin-independent term $(1-\eta){\cal V}_{s}(z)\sigma_{0}$ is mainly contributed from the scalar components of both the $x$- and $y$-polarizations.
@@ VI Conclusion and outlook
We have proposed a generic spin-$1/2$ quasiperiodic (QP) system that unifies the existing important 1D QP models, some of which requiring an extra Majorana representation, and established a unified framework for all fundamental localization phases in quasiperiodic systems. The framework is built on three universal results. First, we showed with renormalization group method the criteria for obtaining the pure phases, namely when the system has no mobility edges. Second, we uncovered a new universal mechanism for the emergence of critical states, as characterized by the generalized incommensurate zeros in matrix elements that remain invariant under dual transformation and introduce critical orbitals into the system. Third, we identified the condition for the exact solvability of QP systems from local constraint, which also provides the mechanism to realize effective unbounded on-site potential in spinful QP chains. These results establish a profound unified theory for all fundamental localization phases in quasiperiodic systems, and offer the key insights for constructing their exactly solvable models. In particular, we proposed the novel spin-selective QP lattice model and QP optical Raman lattice model, in which we predict all basic types of mobility edges (MEs) and all the seven fundamental phases in Anderson localization physics, respectively, which have not been achieved anywhere else. Finally, the experimentally feasible schemes to realize the theoretical predictions have been proposed and studied in detail.
This work enables a broad exploration and exact characterization of the fundamental quantum phases in Anderson localization and opens up intriguing topics for the future study. For example, while the present study is focused on the spin-$1/2$ QP systems, it is natural to extend the current study to the QP systems with larger spins, e.g. to the 1D $\mathrm{SU}(N)$ QP chain with $N$ larger than $2$. Generalizing the mechanism of IDZs in matrix elements for critical states to the $\mathrm{SU}(N)$ systems is of great interests to obtain the rigorous critical states in generic QP systems. The exact characterization of the distinct quantum phases in such $\mathrm{SU}(N)$ QP system could potentially be connected to the exactly solvable QP mosaic models (Wang et al., 2020a; Zhou et al., 2023) with large unit cells, and deserves particular future efforts. On the other hand, extending the present results to higher dimensions is another important issue for the future study.
@@ Acknowledgements.
This work was supported by National Key Research and Development Program of China (Grants No. 2021YFA1400900, No. 2022YFA1405800 and No. 2020YFA0713300), the National Natural Science Foundation of China (Grants No. 12425401, No. 12261160368 and No. 12401208), the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302000), the Shanghai Municipal Science and Technology Major Project (Grant No. 2019SHZDZX01), the Natural Science Foundation of Jiangsu Province (Grant No. BK20241431), the Nankai Zhide Foundation, and by the Fundamental Research Funds for the Central Universities (Grant No. 100-63253099).
@@
Appendix B Proof of criteria for the pure phases without MEs
@@
Under the commensurate approximation of the irrational parameter, the system displays periodic structure and exhibit band dispersion. And the characteristic polynomial is given by the determinant $P^{(n)}(E;\kappa_{x},\kappa_{y})=|{\cal H}^{(n)}-E|$, which can be expanded over the principle dispersion
@@
$+\epsilon_{R}^{(n)}(E,\varphi,\kappa)+T_{R}^{(n)}(E).$ (68)
The system exhibits pure phases without MEs correspond to the case that, the coefficients associated with the relevant dispersions are energy independent. In the presence of the commensurate approximation, then the band dispersion is periodic with respect to both $\kappa_{x}=Lk_{x}$ and $\kappa_{y}=Lk_{y}$. Here, $k_{x}$ is the twisted momentum attached to the hopping coupling matrix $\Pi_{j}\rightarrow\Pi_{j}e^{ik_{x}}$ and $k_{y}$ is the phase offset in the QP modulation $V_{j}^{\mathrm{d}}$ and $V_{j}^{\mathrm{od}}$. Since $P(E)$ has only real roots as $H$ is Hermitian, so the dispersion and $P(E;\kappa_{x},\kappa_{y})$ shares similar periodicity, thus $P(E;\varphi,\kappa)$ is also periodic with respect to $\kappa_{x}$ and $\kappa_{y}$.
Before proving the absence of the MEs when satisfying the criteria in the main text, we first establish the condition that enforces chiral symmetry. In the presence of chiral symmetry, one readily shows that the Hamiltonian excludes any spin-independent, i.e. $\sigma_{0}$ term. We then diagonalize the on-site matrix $M_{j}$ as $M_{j}=m_{j}^{z}\sigma_{z}$, and the resulted hopping coupling matrix retains the form $\Pi_{j}=\sum_{s}p_{j}^{s}\sigma_{s}$, with $s$ being $s=\{x,y,z\}$. When the hopping coupling matrix satisfy
@@
then the system exhibits the chiral symmetry. We simplify notation by setting $m_{j}^{z}\equiv m_{j}$. And Due to the condition in Eq. (69), the transverse hopping coupling components can be parameterized as
@@
with $p_{j}^{\perp}=\sqrt{|p_{j}^{x}|^{2}+|p_{j}^{y}|^{2}}$ and $\tan\theta=p_{j}^{x}/p_{j}^{y}$. Then the system can be rewritten in a more compact from by a unitary transformation
$U(z,\theta)=\left(\begin{array}[]{cc}e^{i\theta}&0\\ 0&e^{-i\theta}\end{array}\right),$ (71)
then we have the eigenvalue equation
@@
with the transformed hopping coupling matrix $\Pi^{\prime}_{j}=U(z,\theta/2)\Pi_{j}U^{\dagger}(z,\theta/2)$ being
@@
and $\vec{v}_{j}=U(z,\theta)\vec{u}_{j}=(v_{j\uparrow},v_{j\downarrow})^{\intercal}$. Thus, under the unitary operator ${\cal O}=\sigma_{y}$, both coupling matrices anticommute
@@
Therefore both $\vec{v}_{j}$ and ${\cal O}\vec{v}_{j}$ are the eigenstate of the system, with the corresponding energies come in pair as $(E,-E)$, confirming the chiral symmetry.
With this properties, we proceed to calculate the characteristic polynomial $P(E)$. We first consider the eigen-equation $H\lvert v\rangle=E\lvert v\rangle$ under the twisted boundary condition
$\begin{array}[]{c}e^{ik_{x}}(p_{j-1}^{z}v_{j-1\uparrow}+p_{j-1}^{\perp}v_{j-1\downarrow})+e^{-ik_{x}}(p_{j}^{z*}v_{j+1\uparrow}+p_{j}^{\perp*}v_{j+1\downarrow})+m_{j}v_{j\uparrow}=Ev_{j\uparrow},\\ e^{ik_{x}}(p_{j-1}^{\perp}v_{j-1\uparrow}-p_{j-1}^{z}v_{j-1\downarrow})+e^{-ik_{x}}(p_{j}^{\perp*}v_{j+1\uparrow}-p_{j}^{z*}v_{j+1\downarrow})-m_{j}v_{j\downarrow}=Ev_{j\downarrow},\end{array}$ (75)
then we locally rotate the bases as $\psi_{j}^{\pm}=v_{j\uparrow}\pm iv_{j\downarrow}$, and we can rewrite the eigen-equation as
@@
By the commensurate approximation, we approximate the irrational number $\alpha$ by the commensurate size, then $L$ becomes the size of the unit cell and the system exhibits the periodic structure. Then to calculate the characteristic polynomial $P(E)=\det|H-EI|$, we can perform the Fourier transformation, i.e. we multiply by $e^{i2\pi\alpha jm}$ and sum over $j$, then finding characteristic polynomial $P(E)=\det|H-EI|$ becomes finding the determinant $P(E)=\det|A(E)|$, with $A(E\text{)}$ being
$A=\left(\begin{array}[]{cccc}A_{11}&A_{12}&\dots&A_{L1}\\ A_{21}&\ddots&&\vdots\\ \vdots&&\ddots&\vdots\\ A_{L1}&\dots&\dots&A_{LL}\end{array}\right)_{2L\times 2L},$
by denoting the hopping coefficients $t_{j}^{\pm}=p_{j}^{z}\pm ip_{j}^{\perp}$, each block being a 2-by-2 matrix as
@@
As mentioned in the main text, this structure ensures the energy terms and momentum dispersions decouple naturally, eliminating cross terms such as $Ee^{ik_{x}}$. Our task is to separately show that the dominant dispersions for extended, localized, and critical states have energy-independent coefficients.
@@
Then we consider the case both $\Pi_{j}$ and $M_{j}$ are purely quasiperiodic, in which case the leading dispersion is $\cos 2Lk_{x}\cos Lk_{y}$, $\cos Lk_{x}\cos 2Lk_{y}$ or $\cos 2Lk_{x}\cos 2Lk_{y}$, whose coefficients are all energy independent. A key ingredient is that the hopping coupling matrix exhibits the off-diagonal modulation $(V_{j}^{\mathrm{od}}c_{j+1s}^{\dagger}c_{js^{\prime}}+\mathrm{h.c.})$, which makes the $p_{j}$ contributes the $k_{x}$ and $k_{y}$ simultaneously. Therefore, if $\Pi_{j}$ are quasiperiodic, $\cos Lk_{x}\cos Lk_{y}$ dispersion will disappear. The reason is that $\cos Lk_{x}\cos Lk_{y}$ comes from the multiplication of $L$ folds $p_{j}$ terms and $L$ folds $E$ terms, which is not allowed under the chiral symmetry. So the leading order contributing to the critical orbital in this case is $\cos Lk_{x}\cos 2Lk_{y}$ or $\cos 2Lk_{x}\cos Lk_{y}$. The first case corresponds multiplication of $L$ folds $p_{j}$ terms and $L$ folds $m_{j}$ terms, which are in total $2L$ terms and leaves no chance to involve the extra terms. So the coefficients associated with $\cos Lk_{x}\cos Lk_{y}$ are energy independent. Similarly, the $\cos 2Lk_{x}\cos Lk_{y}$ originates from the product of $2L$ folds $p_{j}$ terms. Generically this will lead to $\cos 2Lk_{x}\cos Lk_{y}$ and $\cos 2Lk_{x}\cos 2Lk_{y}$, and both of them are energy independent. It worthwhile to note that he former one comes from the destruction of relative phases, which gives rise to the $\cos 2Lk_{x}\cos Lk_{y}$.
Appendix C Generalized 2D bilayer mapping

### v5

@@ IV.3 Unified picture for the seven fundamental localization phases
Reducing the chiral parameter $\eta$ from 1 to 1/2 corresponds to turning on the on-site potential $V_{0}$ in the type-II QP mosaic model, transitioning from $V_{0}=0$ to $V_{0}=t$. This breaks the chiral symmetry of both models and introduces the MEs between localized and critical states, leading to the L+C phase. As the on-site potential increases towards the high-symmetry line $V_{0}=t$, or when the QP optical Raman lattice model reaches $\eta=1/2$, the system becomes analytically solvable again. The type-II QP mosaic model exhibits analytic MEs $E=\pm\lambda$ separating the localized and critical states Zhou et al. (2023), which corresponds to the MEs $E=\pm 2t_{0}$ in the QP optical Raman lattice model.
@@ V Schemes for experimental realization
Finally, we propose experimental schemes for realizing the SSQP on-site model [Eq. (31)] and QP optical Raman lattice model [Eq. (35)]. These schemes utilize ultracold alkali atoms, which are ideal candidates due to their large fine-structure energy splitting and moderate natural linewidth. These properties facilitate efficient Raman coupling and spin-dependent potentials, even when the laser frequency is far detuned from the $D_{1}$ and $D_{2}$ lines Soltan-Panahi et al. (2011); Kuzmenko et al. (2019); Szulim et al. (2022), making ultracold alkali atoms well-suited for implementing the QP lattice. We further outline experimentally accessible observables for detecting the distinct phases and the associated mobility edges.
@@
The proposed Hamiltonian in Eq. (52) can be realized based on QP optical Raman lattice, as outlined below. The detailed implementation can be found in Supplement supplement . To facilitate the description, we perform a unitary transformation, such that $\sigma_{z}\rightarrow\sigma_{x}$ and $\sigma_{x}\rightarrow-\sigma_{z}$. The kinetic term and spin-independent potential ${\cal V}_{s}(z)$ remain unchanged and can be generated using the standing wave ${\bf E}_{1}$. The rotated spin-dependent potential ${\cal V}_{p}(z)(\sigma_{0}-\sigma_{x})/2$ can be realized using a standing waves ${\bf E}_{2}$ and ${\bf E}_{3}$ along with a traveling wave ${\bf E}_{4}$. Finally, the $M_{0}\sigma_{z}$ term arises from the two-photon detuning of the Raman coupling. After performing the inverse unitary transformation, we reach the Hamiltonian in Eq. (52). More details can be found in supplement supplement , where $^{87}{\rm Rb}$ atoms are used as an example.
@@
The different phases can be distinguished by the dynamical exponent $\alpha$ governing wave-packet spreading, $\sigma(t)\sim t^{\alpha}$, where $\sigma(t)$ denotes wave packet width. The exponent takes $\alpha=1$ for extended phase, $0<\alpha<1$ for critical phase, and $\alpha=0$ for localized phase. By preparing initial states at different energies and measuring the corresponding expansion dynamics, one can observe energy-dependent (energy-independent) dynamical exponents, providing direct experimental evidence for mobility edges (pure phases).
@@ VI Conclusion and outlook
We have proposed a broad class of spin-$1/2$ quasiperiodic (QP) system that unifies the existing important 1D QP models, some of which requiring an extra Majorana representation, and established a unified framework for all fundamental localization phases in quasiperiodic systems. The framework is built on three universal results. First, we showed with renormalization group method the criteria for obtaining the pure phases: the mobility edges are absent when the system has chiral(-like) symmetry. Second, we uncovered a new fundamental mechanism for the emergence of critical states, as characterized by the generalized incommensurate zeros in matrix elements that remain invariant under dual transformation and introduce critical orbitals into the system. Third, we identified the condition for the exact solvability of QP models from local constraint, which also provides the mechanism to realize effective unbounded on-site potential in spinful QP chains. These results establish a profound unified theory for all fundamental localization phases in quasiperiodic systems, and offer the key insights for constructing their exactly solvable models. In particular, we proposed the novel spin-selective QP lattice model and QP optical Raman lattice model, in which we predict all basic types of mobility edges (MEs) and all the seven fundamental phases of localization physics in QP systems, respectively, which have not been achieved anywhere else. The experimentally feasible schemes for realization have been proposed and studied in detail.
This work enables a broad exploration and exact characterization of the fundamental quantum phases in Anderson localization and opens up intriguing topics for the future study. For example, while the present study is focused on the spin-$1/2$ QP systems, it is natural to extend the current study to the QP systems with larger spins, e.g. to the 1D $\mathrm{SU}(N)$ QP chain with NN larger than $2$. Generalizing the mechanism of IDZs in matrix elements for critical states to the $\mathrm{SU}(N)$ systems is of great interests to obtain the rigorous critical states in generic QP systems. The exact characterization of the distinct quantum phases in such $\mathrm{SU}(N)$ QP system could potentially be connected to the exactly solvable QP mosaic models (Wang et al., 2020a; Zhou et al., 2023) with large unit cells, and deserves particular future efforts. On the other hand, extending the present rigorous results to the system with long-range hopping, ladder geometries, and further to higher dimensions is another important issue for the future study.
@@ Acknowledgements.
This work was supported by National Key Research and Development Program of China (No. 2021YFA1400900, No. 2022YFA1405800 and No. 2020YFA0713300), the National Natural Science Foundation of China (No. 12425401, No. 12261160368, No. 12401208, No. 12531006 and No. 12526201), the Innovation Program for Quantum Science and Technology (No. 2021ZD0302000), the Shanghai Municipal Science and Technology Major Project (No. 2019SHZDZX01), the Natural Science Foundation of Jiangsu Province (No. BK20241431), the Nankai Zhide Foundation, and by the Fundamental Research Funds for the Central Universities (Grant No. 100-63253099).
@@
Appendix B 1D QP mosaic model
In this section, we introduce the one-dimensional QP mosaic model Wang et al. (2020a); Zhou et al. (2023) and rewrite it into our 1D spinful QP framework. The generic QP mosaic lattice refers that the modulation of on-site potential and hopping coefficients are in a mosaic manner, namely the the QP modulation or the uniform modulation only appears every $\kappa$ sites, as we will see in the following.
where the on-site potential is given by
$V_{j}=\begin{cases}V_{0}\cos(2\pi\alpha j+\theta),&j=0\quad{\rm mod}\penalty 10000\ \kappa,\\ 0,&j\neq 0\quad{\rm mod}\penalty 10000\ \kappa.\end{cases}$ (59)
and the on-site potential given by
Extended Critical Localized
Type I $E<|\lambda^{2}/V_{0}|$ none $E>|\lambda^{2}/V_{0}|$
Type II none $E<|\lambda|$ $E>|\lambda|$
$V_{j}^{\mathrm{M,II}}$ $=$ $2tV_{j}^{\mathrm{d}}(\sigma_{0}+\sigma_{x}).$ (65)
Appendix C Proof of criteria for the pure phases without MEs
@@
Under the commensurate approximation of the irrational parameter $\alpha_{n}=p_{n}/q_{n}$, with the size of rational unit cell fixed to $L=q_{n}$, the system displays periodic structure and band dispersion. And the band dispersion is periodic with respect to both $\kappa_{x}=Lk_{x}$ and $\kappa_{y}=Lk_{y}$. Here, $k_{x}$ is the twisted momentum attached to the hopping coupling matrix $\Pi_{j}\rightarrow\Pi_{j}e^{ik_{x}}$ and $k_{y}$ is the phase offset in the QP modulation $V_{j}^{\mathrm{d}}$ and $V_{j}^{\mathrm{od}}$. The characteristic polynomial is given by the determinant $P^{(n)}(E;\kappa_{x},\kappa_{y})=|{\cal H}^{(n)}-E|$, which can be expanded over the principle dispersion
@@
$+\epsilon_{R}^{(n)}(E,\kappa_{x},\kappa_{y})+T_{R}^{(n)}(E).$ (66)
where $t_{R}^{(n)}(E)$, $V_{R}^{(n)}(E)$ and $\mu_{R}^{(n)}(E)$ are renormalized coefficients of fundamental dispersion $\cos(\kappa_{x}+\kappa_{x}^{0})$, $\cos(\kappa_{y}+\kappa_{y}^{0})$ and $\cos(\kappa_{x}+\tilde{\kappa}_{x}^{0})\cos(\kappa_{y}+\tilde{\kappa}_{y}^{0})$, respectively. The $\kappa_{x}^{0}$, $\kappa_{y}^{0}$, $\tilde{\kappa}_{x}^{0}$ and $\tilde{\kappa}_{y}^{0}$ are model-dependent phase offsets. The term $\epsilon_{R}^{(n)}$ collects the renormalized contributions from higher harmonic (higher-frequency) modes, while $T_{R}^{(n)}(E)$ denotes the dispersionless renormalized coupling independent of $\kappa_{x}$ and $\kappa_{y}$.
In the presence of chiral(-like) symmetry, the Hamiltonian excludes spin-independent ($\sigma_{0}$) term. We then diagonalize the on-site matrix $M_{j}$ as $M_{j}=m_{j}^{z}\sigma_{z}$, and the resulted hopping coupling matrix retains the form $\Pi_{j}=\sum_{s}p_{j}^{s}\sigma_{s}$. When hopping coupling matrices satisfy
@@
then the system exhibits the chiral symmetry. We simplify notation by setting $m_{j}^{z}\equiv m_{j}$. The transverse hopping coupling components can be parameterized as
@@
with $p_{j}^{\perp}=\sqrt{|p_{j}^{x}|^{2}+|p_{j}^{y}|^{2}}$ and $\tan\theta=p_{j}^{x}/p_{j}^{y}$. Under a unitary transformation $U(z,\theta)={\rm diag}(e^{i\theta},e^{-i\theta})$, the eigenvalue equation can be rewritten as
@@
with the transformed hopping coupling matrix being
@@
thus, under the unitary operator ${\cal O}=\sigma_{y}$, both coupling matrices anticommute
@@
Therefore both $\vec{v}_{j}$ and ${\cal O}\vec{v}_{j}$ are the eigenstate of the system, with the corresponding energies come in pair as (E,E)(E,-E), confirming the chiral(-like) symmetry.
With this properties, we proceed to calculate the characteristic polynomial $P(E)$. The eigen-equation $H\lvert v\rangle=E\lvert v\rangle$ under the rotation $\psi_{j}^{\pm}=v_{j\uparrow}\pm iv_{j\downarrow}$ is given by
@@
We further perform the Fourier transformation, then $P(E)=\det|H-EI|=\det|(A_{mj})(E)_{L\times L}|$. By denoting the hopping coefficients $t_{j}^{\pm}=p_{j}^{z}\pm ip_{j}^{\perp}$, each block being a 2-by-2 matrix is given by
@@
In the following, we show that the dominant dispersions have energy-independent coefficients.
@@
To formalize this argument, we analyze the determinant structure of the Fourier-transformed matrix $A(E)$. For uniform $\Pi_{j}$, the leading momentum-dependent contributions to $\det A$ can be systematically extracted by expanding the determinant over permutations and retaining only the lowest-frequency harmonics.
When $\Pi_{j}$ is uniform, the leading coefficients can be obtained explicitly by isolating the dominant contributions in the permutation expansion of $\det A$. In general,
$\det A=\sum_{\sigma\in S_{2L}}\mathrm{sgn}(\sigma)\prod_{m,s}A_{ms;\sigma_{ms}},$ (74)
where $S_{2L}$ is the permutation group of of indexes $1,\ldots,2L$, and $s=\pm 1$ labels internal indices within each block $A_{mj}$. The relevant coefficients arise from contributions of the form
$\sum_{\sigma}\mathrm{sgn}(\sigma)\prod_{m,s}\exp\!\left(i2\pi\alpha ms\sigma_{ms}\right),$ (75)
multiplied by the following momentum-dependent factors
$\Big\{$ $\Big[\Big(\prod_{m}t^{+}e^{-i(2\pi\alpha m+k_{x})}\Big)\Big(\prod_{j}m^{z}\cos(2\pi\alpha j+k_{y})\Big)+\mathrm{h.c.}\Big]$
$+(t^{+}\leftrightarrow t^{-*})$
$+\Big[\Big(\prod_{m}t^{+}e^{-i(2\pi\alpha m+k_{x})}\Big)\Big(\prod_{m}t^{-}e^{-i(2\pi\alpha m+k_{x})}\Big)+\mathrm{h.c.}\Big]$
$+\Big[\prod_{j}m^{z}\cos(2\pi\alpha j+k_{y})\Big]^{2}\Big\}.$ (76)
Retaining only the lowest-frequency harmonics, the determinant reduces to
$\det A\simeq$
$\xi_{0}\Big[$ $\xi_{C}\Big(\frac{m^{z}}{2}\Big)^{L}(|t^{+}|^{L}+|t^{-}|^{L})\cos(Lk_{x}+\tilde{\kappa}_{x}^{0})\cos(Lk_{y})$
$+$ $\xi_{T}|t^{+}t^{-}|^{L}\cos(2Lk_{x}+\kappa_{x}^{0})+\xi_{V}\Big(\frac{m^{z}}{2}\Big)^{2L}\cos(2Lk_{y})\Big]$
$+$ $\text{(higher-frequency terms)}.$ (77)
Here the common prefactor $\xi_{0}$ is given by
$\xi_{0}=\sum_{\sigma}\mathrm{sgn}(\sigma)\prod_{m,s}\exp\!\left(i2\pi\alpha ms\sigma_{ms}\right),$ (78)
and plays no role in the scaling analysis. The coefficients $\xi_{C}$, $\xi_{T}$, and $\xi_{V}$ are nonuniversal constants of order unity, while the phases $\kappa_{x}^{0}$ and $\tilde{\kappa}_{x}^{0}$ depend solely on $\arg(t^{\pm})$. From this expression, the effective renormalized coefficients for the leading dispersions can be directly obtained
$t_{2R}^{(L)}$ $\approx|t^{+}t^{-}|^{L},$ (79)
$V_{2R}^{(L)}$ $\approx\Big(\frac{m^{z}}{2}\Big)^{2L},$ (80)
$C_{R}^{(L)}$ $\approx\Big(\frac{m^{z}}{2}\Big)^{L}\big(|t^{+}|^{L}+|t^{-}|^{L}\big).$ (81)
These results align with phase boundary in Table. 2 Wang et al. (2020b); Gonçalves et al. (2023b). In this derivation, we assume that intermediate power combinations, such as $(t^{+})^{L-1}(t^{-})^{L+1}$, do not contribute at leading order. Such terms enter the permutation sum with effectively random phases and cancel out, leading to a parametric suppression relative to the coherent lowest-harmonic terms.
Then we consider the case both $\Pi_{j}$ and $M_{j}$ are purely QP, in which case the leading dispersion is $\cos Lk_{x}\cos 2Lk_{y}$ or $\cos 2Lk_{x}\cos 2Lk_{y}$, whose coefficients are all energy independent. A key ingredient is that the hopping coupling matrix exhibits the off-diagonal modulation $(V_{j}^{\mathrm{od}}c_{j+1s}^{\dagger}c_{js^{\prime}}+\mathrm{h.c.})$, which makes the $p_{j}$ contributes the $k_{x}$ and $k_{y}$ simultaneously. Therefore, if $\Pi_{j}$ are QP, $\cos Lk_{x}\cos Lk_{y}$ dispersion will disappear. The reason is that $\cos Lk_{x}\cos Lk_{y}$ comes from the multiplication of $L$ folds $p_{j}$ terms and $L$ folds $E$ terms, which is not allowed under the chiral(-like) symmetry. So the leading order contributing to the critical orbital in this case is $\cos Lk_{x}\cos 2Lk_{y}$ or $\cos 2Lk_{x}\cos Lk_{y}$. The first case corresponds multiplication of $L$ folds $p_{j}$ terms and $L$ folds $m_{j}$ terms, which are in total $2L$ terms and leaves no chance to involve the extra terms. And the latter dispersion cannot be realized in the system. So the coefficient associated with $\cos Lk_{x}\cos 2Lk_{y}$ is energy independent. Similarly, the $\cos 2Lk_{x}\cos 2Lk_{y}$ originates from the product of $2L$ folds $p_{j}$ terms and is also energy independent.
Appendix D Generalized 2D bilayer mapping